1) Explain the meaning of absolute zero on the Kelvin scale:

a. is the freezing point of water

b. gas molecules quit moving at this temperature

c. is a practical limit, theory allows lower

d. heat can flow from an object at this temperature

2) What characteristic of water makes it ideal as a practical engine coolant?

a. highest specific heat

b. lowest specific heat

c, highest boiling point

d. lowest freezing point

3) A piston has an area of 3×10−2 m2 and a force on it of 7.5×105 N. What is its pressure in Pa?

a. 225

b. 250

c. 2.25×107

d. 2.5×107

4) How much heat is needed to raise 50 g of aluminum from 20 °C to 100 °C?

a. 110 cal

b. 1200 cal

c. 860 cal

d. 1500 cal

5) 50 g of water at 20°C is mixed with 80 g of water at 60 °C. What Is the mixture’s final temperature?

a. 44.6 °C

b. 83.2 °C

c. 95.2 °C

d. 25.3 °C

6) A negatively charged body has

a. Lost electrons

b. Gained electrons

c. Lost protons

d. Gained protons
Answer #1

2.5X10pa Given A= 3x10² m² F= 7.580N pressure is given as - p=t 7.58,05N 3x152 m2 P=2.5x10tPa 6 is correct option. 4 Given,50 mic, DT, = M₂ CDTz 620-1) = 80 gw (7-80) choice 20-T = 89 (7-8 > 20-T = 8 T-9 T($+1) = 96+20 T (1) = 116 11645 ! Ta - 44.6

Hope this helps you.

Please thumbs up.